# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

### v1

@@ VII-A Experiment Setup
Transition Model Model learning rate $1\times e^{-3}$
@@
Reward Model Model learning rate $3\times e^{-4}$
@@
Policy Model Actor learning rate (IQL) $1(3)\times e^{-4}$
Critic learning rate $3\times e^{-4}$
@@
Experimental Details. Offline PbRL setting involves two steps: reward learning and policy learning. Here, we directly use offline RL algorithms CQL [4] and IQL [25] to perform policy learning. The performance of these algorithms is evaluated based on the cumulative return. For comparison, the scores are normalized between 0 (random policy score) and 100 (expert policy score) [22]. The normalized score $\tilde{J}$ is computed by:
$\tilde{J}=\frac{J-J_{r}}{J_{e}-J_{r}}\times 100,$ (16)
where $J_{r}$, $J_{e}$ and $J$ are the expected return of a random policy, an expert policy, and the trained policy by offline RL algorithms, respectively. Table I gives the hyper-parameter configuration for reward and transition training, and policy learning. The details are given the below:
Figure 3: The relationship between preference model accuracy and the number of preference dataset $N_{l}$.
TABLE II: The hyperparameters for LEASE under CQL and IQL algorithms.
Task Name CQL IQL
$H$ $\kappa_{p}$ $\kappa_{\tau}$ $H$ $\kappa_{p}$ $\kappa_{\tau}$
walker2d-m $10$ $0.85$ $0.05$ $10$ $0.75$ $0.08$
walker2d-m-e $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
hopper-m $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
hopper-m-e $100$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
halfcheetah-m $10$ $0.85$ $0.05$ $10$ $0.85$ $0.08$
halfcheetah-m-e $200$ $0.85$ $0.05$ $10$ $0.75$ $0.08$
pen-human $10$ $0.95$ $0.05$ $10$ $0.85$ $0.08$
pen-expert $10$ $0.90$ $0.05$ $10$ $0.95$ $0.08$
door-human $10$ $0.95$ $0.05$ $10$ $0.90$ $0.08$
door-expert $20$ $0.95$ $0.05$ $10$ $0.95$ $0.08$
hammer-human $10$ $0.95$ $0.05$ $10$ $0.90$ $0.08$
hammer-expert $20$ $0.99$ $0.05$ $10$ $0.95$ $0.08$
1) Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models. Each model consists of a $4$-layer feed-forward neural network with $200$ hidden units. The model training employs maximum likelihood estimation with a learning rate of $1\times e^{-3}$ and Adam optimizer. Following [21], we train an ensemble of $3$ reward models. Each reward model includes a $3$-layer feed-forward neural network with $256$ hidden units. The reward model training uses cross-entropy loss with a learning rate of $3\times e^{-4}$ and Adam optimizer.
2) Policy optimization: The offline algorithms include CQL and IQL. In CQL algorithm, the actor network $\pi_{\theta}$ and critic network $Q_{\omega}$ adopt a $3$-layer feed-forward neural network with learning rate of $1\times e^{-4}$ and $3\times e^{-4}$, respectively. The cql weight and temperature are set as $5.0$ and $1.0$ for all tasks, respectively. In IQL algorithm, the actor network and critic network adopt a $2$-layer neural network with learning rate of $3\times e^{-4}$ and $3\times e^{-4}$, respectively. The expectile and temperature are set as $0.7$ and $3.0$ for mujoco tasks and $0.8$ and $3.0$ for adroit tasks.
3) Parameter tuning: Theorem 1 indicates that the performance of reward model is influenced by the selecting mechanism $f(\sigma_{0}^{u},\sigma_{1}^{u})$ and the quality of the generated data. Therefore, the key hyperparameters should be tuned for LEASE including: rollout horizon $H\in\{10,20,100,200\}$, probability confidence $\kappa_{p}(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})\in\{0.75,0.85,0.90,0.95,0.99\}$ and uncertainty variance $\kappa_{\tau}(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})\in\{0.05,0.08\}$. Table II shows the above hyperparameters for LEASE under CQL and IQL algorithms.
@@
VII-B Results for Reward Model
The effect of the number of preference data for preference model. Learning an accurate reward is difficult in the offline PbRL setting when preference labels are scarce. Fig. 3 illustrates the relationship between the prediction accuracy of preference model without data augmentation and the number of preference data $N_{l}$, where preference model is based on reward model (Eq. (2)). Here, we use the last $200$ of $2000$ preference data that are not seen in training stage as evaluation dataset. It indicates that as the number of preference data $N_{l}$ increases, the prediction accuracy of the preference model shows an upward trend. This also validates Theorem 1 that increasing the number of samples $N_{l}$ can reduce the generalization error.
@@ VII-B Results for Reward Model
TABLE III: Comparison results of pseudo-labeling accuracy and agent performance between using confidence and uncertainty and only using confidence. The accuracy of pseudo-label generated by reward model is evaluated on all preference dataset.
Selecting Mechanism pen-expert door-expert hammer-expert
label accuracy agent performance label accuracy agent performance label accuracy agent performance
Confidence and uncertainty $87.3\pm 0.5$ $\mathbf{132.5\pm 2.3}$ $89.3\pm 0.6$ $\mathbf{103.2\pm 0.7}$ $85.5\pm 0.2$ $\mathbf{126.3\pm 1.2}$
Only confidence $85.9\pm 0.8$ $126.8\pm 0.1$ $87.8\pm 0.2$ $101.1\pm 2.2$ $84.4\pm 0.4$ $118.2\pm 9.6$
The effect of data augmentation for the performance of reward model. We compare the accuracy of reward model before and after data augmentation under fewer preference dataset. Fig. 4 shows the comparison results between prediction and ground truth of reward, where the predicted and true rewards are both normalized to $[0,1]$. We randomly sample $500$ data from unlabeled datasets that are not seen in training stage for evaluation. This figure indicates the linear relationship between reward predicted by LEASE and ground truth is better than that of other two methods, where method not using selecting mechanism is denoted as FRESH. The prediction of reward model learned by method FRESH is very narrow and the accuracy is greatly reduced compared with FEWER. This is because generated data of FRESH are not selected by $f(\sigma_{0},\sigma_{1})$, which causes substantial errors for the labels of generated data and leads to the collapse of training. This validates that the performance of reward model can be improved through data augmentation and selecting mechanism, where the designed selecting mechanism has a greater impact on reward model performance.
TABLE IV: The performance of the learned transition model, where error is calculated by the sum of the mean square values of predicted and true value in each dimension.
Task Name walker2d-m-e hopper-m-e halfcheetah-m-e
Error $6.57\pm 0.10$ $0.38\pm 0.01$ $16.19\pm 0.48$
Task Name pen-expert door-expert hammer-expert
Error $1.05\pm 0.01$ $0.04\pm 0.01$ $0.58\pm 0.03$
The effect of uncertainty for reward model and agent performance. Table III compares the accuracy of pseudo-labels and performance between the proposed method LEASE and the confidence-only method. The latter method is the Surf [3] that only uses confidence for screening data. This table validates that LEASE outperforms the confidence-only approach in pseudo-labels accuracy and agent performance. The performance improved by $4.5\%$, $2.8\%$ and $6.85\%$ for pen-expert, door-expert and hammer-expert, respectively. Additionally, there is a difference between the accuracy of the preference model $P(\sigma_{0}\succ\sigma_{1})$ (pseudo-label accuracy) and the accuracy of the reward model $R(s,a)$ (See Eq. (2)). A decline in the accuracy of the reward model may not necessarily lead to a significant drop in the accuracy of the preference model, as the preference model is related to the summation of reward model over trajectories.
VII-C Results for Transition model
The accuracy of transition model. Table IV shows the accuracy result for the learned transition model, where the accuracy of transition model is measured by the sum of the mean square values of the predicted value and the true value in each dimension. Note that, for walker2d and halfcheetah tasks, the error of the learned transition model is greatly larger than hopper tasks in mujoco environment. This is mainly because the state space of walker2d and hopper is larger than hopper, and the physical model of them are more complex.
@@ VII-C Results for Transition model
TABLE V: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds. $\pm$ captures the standard deviation over seeds. Bold indicates the performance within 2% of the best performing algorithm for each offline algorithm.
@@
walker2d-m $76.0\pm 0.9$ $77.4\pm 0.6$ $\mathbf{78.4\pm 0.9}$ $\mathbf{78.4\pm 0.5}$ $71.8\pm 0.8$ $74.6\pm 1.8$
@@
hopper-m $54.7\pm 3.4$ $55.8\pm 2.8$ $\mathbf{56.5\pm 0.6}$ $50.8\pm 4.2$ $\mathbf{56.7\pm 2.6}$ $\mathbf{56.0\pm 0.5}$
hopper-m-e $\mathbf{57.4\pm 4.9}$ $53.6\pm 0.9$ $56.4\pm 0.8$ $\mathbf{94.3\pm 7.4}$ $56.0\pm 1.4$ $55.9\pm 1.9$
@@
halfcheetah-m-e $\mathbf{62.7\pm 7.1}$ $48.3\pm 0.7$ $53.2\pm 3.1$ $\mathbf{91.0\pm 2.3}$ $59.1\pm 4.9$ $62.4\pm 1.4$
Mujoco Average $\mathbf{64.5\pm 6.5}$ $59.4\pm 0.9$ $\mathbf{64.4\pm 4.0}$ $\mathbf{77.9\pm 2.5}$ $65.2\pm 2.2$ $66.7\pm 1.0$
@@
pen-expert $\mathbf{138.3\pm 5.2}$ $128.1\pm 0.7$ $132.5\pm 2.3$ $\mathbf{132.9\pm 4.6}$ $104.1\pm 12.9$ $113.8\pm 6.3$
door-human $\mathbf{4.7\pm 5.9}$ $0.2\pm 1.0$ $\mathbf{4.7\pm 8.8}$ $3.5\pm 3.2$ $4.0\pm 2.5$ $\mathbf{5.9\pm 0.5}$
@@
hammer-human $\mathbf{0.9\pm 0.3}$ $0.3\pm 0.0$ $0.3\pm 0.0$ $1.4\pm 1.0$ $1.2\pm 0.2$ $\mathbf{1.7\pm 0.4}$
@@
Adroit Average $\mathbf{63.0\pm 5.5}$ $59.4\pm 1.3$ $61.8\pm 3.0$ $\mathbf{70.1\pm 4.2}$ $67.1\pm 4.8$ $\mathbf{71.4\pm 1.8}$
TABLE VI: The comparison results between the performance using selecting mechanism and that not using. The latter method is denoted as FRESH. $\uparrow$ denotes the improvement of performance.
@@
LEASE $\mathbf{74.6\uparrow(2.6\%)}$ $\mathbf{56.0\uparrow(4.5\%)}$ $\mathbf{43.0\uparrow(1.3\%)}$ $\mathbf{108.1\uparrow(2.7\%)}$ $\mathbf{55.9\uparrow(3.7\%)}$ $\mathbf{62.4\uparrow(14.6\%)}$
TABLE VII: The prediction result of preference model for different rollout length $H$. Pretrained represents the accuracy of preference model trained with fewer preference data, and updated represents that updated with generated data.
Task name walker2d-m-e hopper-m-e halfcheetah-m-e
Pretrained $0.85$ $0.77$ $0.72$
Updated $H=10$ $\mathbf{0.85}$ $0.74$ $\mathbf{0.78}$
$H=100$ $0.84$ $\mathbf{0.79}$ $0.74$
$H=200$ $0.83$ $0.76$ $0.77$
The effect of rollout length $H$ of transition model for preference model. Table VII shows the prediction accuracy of the trained preference model for different rollout length $H$. Here, we take CQL as example and select the last $200$ of $2000$ preference data as evaluation dataset that are not seen in in reward training stage. The accuracy is calculated through the gap between predicted preference and true preference based on two trajectories. It shows that long horizon easily brings low performance of reward model since the cumulative error of the trained transition model, but long horizon is beneficial to enhance the performance of reward model under some conditions. The accuracy of preference model can be improved through data augmentation under most tasks.
Note that the rollout length $H$ is equal to the length of generated unlabeled data $L$. The $H$ is influenced by the accuracy of trained transition model. As $H$ increases, the accuracy of prediction decreases and large error will bring the unstable agent training. However, the long horizon of trajectory can better describe the preference of human feedback and can reduce sample complexity (See Eq. (15)). Additionally, the results in table VII can not completely reflect the generalization performance of reward model since the evaluation data are limited and there is difference between preference model and reward model (See Eq. (2)).
TABLE VIII: Results of the effect of the learned transition model for agent performance.
No. hopper-medium pen-expert
Error Performance Error Performance
$1$ $0.35\pm 0.01$ $56.5\pm 0.60$ $1.05\pm 0.01$ $132.5\pm 2.3$
$2$ $0.49\pm 0.04$ $54.8\pm 0.85$ $1.42\pm 0.05$ $126.4\pm 8.6$
$3$ $1.19\pm 0.05$ $52.9\pm 0.92$ $2.36\pm 0.24$ $87.7\pm 41.8$
The effect of the accuracy of transition model for agent performance. Table VIII shows the detailed results for the effect of transition model accuracy for agent performance, where the smaller the model error means the higher the accuracy. The change in transition model error is achieved by changing the number of training steps. The base offline RL algorithm is CQL. It shows that the lower the accuracy of the transition model, the poorer performance of the agent.
VII-D Results for Policy Performance
The comparison results on benchmark tasks. Here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms. Table V gives the results of the average normalized score with standard deviation, where m and m-e denote medium and medium-expert respectively. The offline PbRL algorithm under sufficient feedback (URLHF) are obtained from paper [21]. This table shows that LEASE can greatly improve performance under fewer preference data compared with FEWER, and achieve comparable performance to URLHF that demand large human feedback. This validates the content of Theorem 2 that LEASE can reduce the performance gap caused by reward model under fewer preference data and has high sample efficiency.
@@ VII-D Results for Policy Performance
TABLE IX: The comparison results for the D4RL tasks under different number of preference data $N_{l}$. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score.
@@
walker2d-medium-expert $98.0\pm 18.6$ $\mathbf{98.6\pm 18.1}$ pen-expert $102.5\pm 12.8$ $\mathbf{113.8\pm 6.3}$
hopper-medium $\mathbf{56.8\pm 1.5}$ ${56.5\pm 0.6}$ door-human $4.4\pm 1.2$ $\mathbf{5.9\pm 0.5}$
@@
halfcheetah-medium $43.4\pm 0.4$ $\mathbf{43.5\pm 0.1}$ hammer-human $1.2\pm 0.4$ $\mathbf{1.7\pm 0.4}$
@@
Mujoco Average $63.6\pm 4.0$ $\mathbf{64.4\pm 4.0}$ Adroit Average $68.6\pm 3.7$ $\mathbf{71.4\pm 1.8}$

### v2

@@ VI-A2 Experimental Details
Transition Model Model learning rate $1\times 10^{-3}$
@@
Reward Model Model learning rate $3\times 10^{-4}$
@@
Policy Model Actor learning rate (IQL) $1(3)\times 10^{-4}$
Critic learning rate $3\times 10^{-4}$
@@
Reward and transition model training: Similar to [37], we train an ensemble of $7$ transition models and select the best $5$ models. Each model consists of a $4$-layer feed-forward neural network with $200$ hidden units. The model training employs maximum likelihood estimation with a learning rate of $1\times 10^{-3}$ and Adam optimizer. Following [21], we train an ensemble of $3$ reward models. Each reward model includes a $3$-layer feed-forward neural network with $256$ hidden units. The reward model training uses cross-entropy loss with a learning rate of $3\times 10^{-4}$ and Adam optimizer.
Policy optimization: The offline algorithms include CQL [4] and IQL [25]. In CQL algorithm, the actor network $\pi_{\theta}$ and critic network $Q_{\omega}$ adopt a $3$-layer feed-forward neural network with learning rate of $1\times 10^{-4}$ and $3\times 10^{-4}$, respectively. The cql weight and temperature are set as $5.0$ and $1.0$ for all tasks, respectively. In the IQL algorithm, the actor network and critic network adopt a $2$-layer neural network with learning rate of $3\times 10^{-4}$ and $3\times 10^{-4}$, respectively. Similar to [25], the expectile and temperature are set as $0.7$ and $3.0$ for mujoco tasks and $0.8$ and $3.0$ for adroit tasks.
[…]
@@
TABLE II: Comparison results of pseudo-labeling accuracy between using the selection mechanism (Ours), only using confidence (No-U), only using uncertainty (No-C), and not using the selection mechanism (No-C-U).
Task Name Ours No-U No-C No-C-U
pen-expert $\mathbf{87.3\pm 0.5}$ $85.9\pm 0.8$ $86.7\pm 0.6$ $86.8\pm 0.5$
door-expert $\mathbf{89.3\pm 0.6}$ $87.8\pm 0.2$ $86.0\pm 0.0$ $85.9\pm 0.8$
hammer-expert $\mathbf{85.5\pm 0.2}$ $84.4\pm 0.4$ $82.5\pm 1.2$ $83.6\pm 0.5$
Average $\mathbf{87.4\pm 0.4}$ $86.0\pm 0.5$} $85.1\pm 0.6$} $85.4\pm 0.6$}
@@
TABLE III: Results for the D4RL tasks during the last 5 iterations of training averaged over 3 seeds. $\pm$ captures the standard deviation over seeds. Bold indicates the performance within 2% of the best-performing algorithm for each offline algorithm.
@@
walker2d-m $6.0\pm 0.9 77$$7.4\pm 0.6$ $\mathbf{78.4\pm 0.9}$ $\mathbf{78.4\pm 0.5}$ $71.8\pm 0.8$ $73.0\pm 2.1$
@@
hopper-m $4.7\pm 3.4 55$$5.8\pm 2.8$ $\mathbf{56.5\pm 0.6}$ $50.8\pm 4.2$ $\mathbf{56.7\pm 2.6}$ $\mathbf{56.0\pm 0.5}$
hopper-m-e $\mathbf{57.4\pm 4.9}$ $53.6\pm 0.9$ $56.4\pm 0.8$ $\mathbf{94.3\pm 7.4}$ $6.0\pm 1.4 55$$5.9\pm 1.9$
@@
halfcheetah-m-e $\mathbf{62.7\pm 7.1}$ $48.3\pm 0.7$ $53.2\pm 3.1$ $\mathbf{91.0\pm 2.3}$ $59.1\pm 4.9$ $60.2\pm 1.2$
Mujoco Average $\mathbf{64.5\pm 6.5}$ $59.4\pm 0.9$ $\mathbf{64.4\pm 4.0}$ $\mathbf{77.9\pm 2.5}$ $5.2\pm 2.2 66$$6.4\pm 1.1$
@@
pen-expert $\mathbf{138.3\pm 5.2}$ $128.1\pm 0.7$ $130.6\pm 1.8$ $\mathbf{132.9\pm 4.6}$ $04.1\pm 12.9 11$$13.8\pm 6.3$
door-human $\mathbf{4.7\pm 5.9}$ $0.2\pm 1.0$ $\mathbf{4.7\pm 8.8}$ $3.5\pm 3.2$ $4.0\pm 2.5$ $\mathbf{5.7\pm 0.8}$
@@
hammer-human $\mathbf{0.9\pm 0.3}$ $0.3\pm 0.0$ $0.3\pm 0.0$ $\mathbf{1.4\pm 1.0}$ $1.2\pm 0.2$ $\mathbf{1.4\pm 0.9}$
@@
Adroit Average $\mathbf{63.0\pm 5.5}$ $59.4\pm 1.3$ $61.5\pm 2.9$ $\mathbf{70.1\pm 4.2}$ $67.1\pm 4.8$ $\mathbf{71.3\pm 1.9}$
VI-B Results for Reward Model
VI-B1 The effect of the selection mechanism for pseudo-label accuracy
We employ a pre-trained reward model to annotate unlabeled preference data generated during the process. This paper proposes a novel screening mechanism to select data samples with both high confidence and low uncertainty, thereby enhancing the accuracy of pseudo-labels. The pseudo-label is generated by a preference model, which is constructed based on the reward model as formulated in Eq. (2). Table II compares the pseudo-label accuracy under four different conditions: 1) using the proposed screening mechanism (combining both confidence and uncertainty thresholds), 2) employing only the confidence threshold (No-U), 3) applying only the uncertainty threshold (No-C), and 4) without any screening mechanism (No-C-U). The accuracy of the pseudo-label generated by the reward model is evaluated on all preference datasets.
Figure 4: The relationship between pseudo-labeling accuracy and the number of preference dataset $N_{l}$.
Table II shows that the proposed screening mechanism significantly improves pseudo-label accuracy, i.e., reduces error $\eta$ defined in Assumption 2. The pseudo-label accuracy is improved by $4.5\%$, $2.8\%$, and $6.9\%$ for pen-expert, door-expert, and hammer-expert, respectively. Moreover, we also analyze the effect of the number of preference data on the preference model. Fig. 4 illustrates the relationship between the prediction accuracy of the preference model without data augmentation and the number of preference data $N_{l}$. Here, we use the last $200$ of $2000$ preference data that are not seen in the training stage as the evaluation dataset. It indicates that as the number of preference data $N_{l}$ increases, the prediction accuracy of the preference model shows an upward trend. This also validates Theorem 1 that increasing the number of samples $N_{l}$ can reduce the generalization error.
VI-B2 The effect of data augmentation on the performance of the reward model
To answer question $\mathbf{Q_{1}}$, we compare the accuracy of the reward model before and after data augmentation under a smaller preference dataset. Fig. 3 shows the comparison results between prediction and ground truth of reward, where the predicted and true rewards are both normalized to $[0,1]$. We randomly sample $500$ data points from unlabeled datasets that are not seen in the training stage for evaluation. This figure indicates that the linear relationship between reward predicted by LEASE and ground truth is better than that of the other two methods, where the method not using a selecting mechanism is denoted as FRESH. The prediction of the reward model learned by method FRESH is very narrow, and the accuracy is greatly reduced compared with FEWER. This is because the generated data of FRESH are not selected by $f(\sigma_{0},\sigma_{1})$, which causes substantial errors for the labels of generated data and leads to the collapse of training. This validates that the performance of the reward model can be improved through data augmentation and a selection mechanism, where the designed selection mechanism has a greater impact on the reward model’s performance.
VI-C Results for Policy Performance
TABLE IV: The comparison results between the performance using the selection mechanism and that not using it. The latter method is denoted as FRESH. $\uparrow$ denotes the improvement of performance.
@@ VI-C Results for Policy Performance
LEASE $\mathbf{73.0\uparrow(0.4\%)}$ $\mathbf{56.0\uparrow(4.5\%)}$ $\mathbf{43.0\uparrow(1.3\%)}$ $\mathbf{108.1\uparrow(2.7\%)}$ $\mathbf{55.9\uparrow(3.7\%)}$ $\mathbf{60.2\uparrow(10.5\%)}$
VI-C1 The comparison results on benchmark tasks
To answer question $\mathbf{Q_{2}}$, here, we evaluate LEASE performance based on CQL [4] and IQL [25] two offline algorithms on six types of tasks and 12 datasets in total. Table III presents a comparative analysis of he results of the average normalized score with standard deviation under three distinct scenarios: 1) performance with extensive preference data (URLHF), 2) performance with limited data without employing data augmentation techniques (FEWER), and 3) performance with limited data while incorporating data augmentation (our proposed method LEASE). In this table, m and m-e represent medium and medium-expert, respectively. The results of the offline PbRL algorithm under sufficient feedback (URLHF) are obtained from the paper [21]. As demonstrated in Table III, LEASE exhibits significantly superior performance compared to FEWER. This empirical evidence substantiates that data augmentation can effectively enhance the agent’s capability.
Furthermore, compared to the performance achieved with extensive preference data URLHF, LEASE attains comparable results in most cases. However, the proposed method exhibits slightly inferior performance on specific tasks, such as hopper-medium-expert and halfcheetah-medium-expert datasets. This limitation primarily stems from the excessively narrow distribution of the expert dataset, which consequently induces a constrained distribution in the generated preference data. As a result, the reward model’s performance cannot be effectively enhanced through data augmentation, thereby restricting the potential improvement in agent performance. However, overall, the results in Table III can validate the content of Theorem 2, which shows that LEASE can reduce the performance gap caused by the reward model and improve the agent’s performance with fewer preference data.
@@ VI-C1 The comparison results on benchmark tasks
TABLE V: The comparison results for the D4RL tasks under different number of preference data $N_{l}$. $\pm$ captures the standard deviation over seeds. Bold indicates the highest score.
@@
[…]
hopper-medium $\mathbf{56.8\pm 1.5}$ $56.5\pm 0.6$} door-human $4.4\pm 1.2$ $\mathbf{5.7\pm 0.8}$
@@
halfcheetah-medium $43.4\pm 0.4$ $\mathbf{43.5\pm 0.1}$ hammer-human $1.2\pm 0.4$ $\mathbf{1.4\pm 0.9}$
@@
Mujoco Average $63.6\pm 4.0$ $\mathbf{64.4\pm 4.0}$ Adroit Average $68.6\pm 3.7$ $\mathbf{71.3\pm 1.9}$
VI-C2 The effect of selection mechanism on agent performance